A mixture of 0. 166 moles of C is reacted with 0.117 moles of O2 in a sealed, 10.0 L vessel at 500.0 K, producing a mixture of CO and CO2 


3C(s)+2O2(g) → 2CO(g) + CO2(g) 


The total pressure is 0.707 atm. What is the partial pressure of CO? 
Answer #1

Step :1 Calculation of limiting reagent.

(Note : Limiting reagent is the compound that completely consumed in the reaction and limits the formation of product).

According to the given reaction

3 moles of carbon reacts with 2 moles of oxygen, then

0.166 moles of carbon reacts with how many moles of oxygen.

Number of moles of oxygen= 0.166 X 2 /3 =0.1106 moles.

But, the given number of moles of oxygen are 0.117 moles, that suggests, oxygen is available in excess than carbon, and hence carbon is the limiting reagent, on which the number of moles of prudupro depends.

Therefore,

Number of moles of un reacted oxygen = Total number of moles of oxygen - number of moles of oxygen consumed.

= 0.117 - 0.1106

= 0.0064 moles.

Step 2: Calculation of number of moles of CO formed.

According to the given reaction

3 moles of carbon can form, 2 moles of CO, then

0.166 moles of carbon can form........

Number of moles of CO = 0.166 X 2/3

= 0.1106 moles.

Step 3: calculation of the number of moles of CO​​​​​2.

As per the given reaction,

3 moles of carbon can form, one mole of Carbon dioxide, then

0.166 moles of carbon can form.........

Number of moles of CO​​​​​2​​​​ = 0.166 X 1/3

= 0.0553 moles.

Step 4 : Calculation of number of moles of gaseous compounds present in the reaction mixture.

Total number of moles of gaseous compounds = number of moles oxygen + number of moles of CO + number of moles CO​​​​​2​​​​.

= 0.0064 + 0.1106 + 0.0553

= 0.1723 moles.

Step 5 : Calculation of the mole fraction of CO

Mole fraction of CO = number of moles of CO / Total number of moles of all the gaseous compounds in the reaction mixture.

Mole fraction of CO = 0.1106/0.1723

= 0.6419.

Step 6 : Calculation of partial pressure of CO.

According to the Dalton's law of partial pressure,

Partial pressure of CO = mole fraction of CO X Total pressure.

P​​​​​​co​​​​ = 0.6419 X 0.707

= 0.4538 atm.